# SpectraLift: Physics-Guided Spectral-Inversion Network for Self-Supervised Hyperspectral Image Super-Resolution

### v1

@@ 4 Ablation
To quantify the impact of our architectural and training choices, we evaluate a suite of SpectraLift/SIN variants on the Washington DC Mall benchmark using only the Gaussian PSF; results are given in Table 11. We perform the experiments using the same procedure as described in Section 3. All variants of SIN share the same hyperparameters and differ only in the component under test. Removing our selective residual links degrades reconstruction quality, confirming that skip connections stabilize the ill-posed spectral inversion by biasing the network when appropriate at almost no additional complexity cost. Disabling the learning-rate scheduler also harms convergence, while swapping the MAE loss for MSE causes a small drop in fidelity. By contrast, replacing MAE with a cosine-similarity loss collapses performance, underlining that we must penalize direct magnitude errors in the spectral domain. Changing activations – Leaky ReLU to ReLU or GeLU – has only a minor effect, showing robustness to the choice of nonlinearity.
SIN depth and width both matter: our chosen six-layer MLP achieves the best balance of accuracy and efficiency. An eight-layer variant offers no additional gains at substantially higher compute, whereas a four-layer or two-layer version remains competitive while having fewer parameters and FLOPs. Just using a MLP with a single hidden layer is also capable or achieving decent quality metrics at substantially lower computational cost. Reducing hidden-unit size further cuts complexity with only slight quality loss, while doubling hidden-unit size does not offer any benefits at much higher model complexity. A purely linear map (single dense layer without nonlinear activation) performs much worse, empirically validating that modeling hyperspectral signals on a near-linear manifold – rather than as a strictly linear mapping – is essential.
@@ 5 Conclusion
We have introduced SpectraLift, a fully self-supervised HSI-SR framework that leverages the well-defined, readily available SRF (or its Gaussian approximations in cases where the exact SRF is not available) to directly invert spectral compression with a compact, per-pixel MLP. By casting fusion as an spectral-inversion problem grounded in a clear physics model, SpectraLift avoids black-box approaches, and delivers interpretable mappings. Extensive evaluations across synthetic and real-world benchmarks show SpectraLift consistently matches or exceeds leading supervised methods on key quality metrics, while outperforming state-of-the-art unsupervised baselines. Its lightweight design and robustness to spatial blur and temporal misalignment further highlight its practical utility. SpectraLift establishes a foundation for lightweight, physics-guided spectral inversion and paves the way for future advances in efficient, transparent HSI-MSI fusion models for HSI-SR.

### v2

@@ 3.4 Ablation
To quantify the impact of our architectural and training choices, we evaluate a suite of SpectraLift/SIN variants on the Washington DC Mall benchmark using only the Gaussian PSF; results are given in Table 11. We perform the experiments using the same procedure as described in Section 3.2. All variants of SIN share the same hyperparameters and differ only in the component under test. Removing our selective skip/residual connections degrades reconstruction quality, confirming that skip connections stabilize the ill-posed spectral inversion by biasing the network when appropriate at almost no additional complexity cost. Disabling the learning-rate scheduler also harms convergence, while swapping the MAE loss for MSE causes a small drop in fidelity. By contrast, replacing MAE with a cosine-similarity loss collapses performance, underlining that we must penalize direct magnitude errors in the spectral domain. Changing activations – Leaky ReLU to ReLU or GeLU – has only a minor effect, showing robustness to the choice of nonlinearity.
SIN depth and width both matter: our chosen six-layer MLP achieves the best balance of accuracy and efficiency. An eight-layer variant offers no additional metric gains at substantially higher compute, whereas a four-layer or two-layer version remains competitive while having fewer parameters and FLOPs. Just using a MLP with a single hidden layer is also capable or achieving decent quality metrics at substantially lower computational cost. Reducing hidden-unit size further cuts complexity with only slight quality loss, while doubling hidden-unit size does not offer any benefits at much higher model complexity. A purely linear map (single dense layer without nonlinear activation) performs much worse, empirically validating that modeling hyperspectral signals on a near-linear manifold – rather than as a strictly linear mapping – is essential.
@@ 4 Conclusion
We have introduced SpectraLift, a fully self-supervised HSI-SR framework that leverages the well-defined, readily available SRF (or its Gaussian approximations in cases where the exact SRF is not available) to directly invert spectral compression with a compact, per-pixel MLP. By casting fusion as an spectral-inversion problem grounded in a clear physics model, SpectraLift avoids black-box approaches, and delivers interpretable mappings. Extensive evaluations across synthetic and real-world benchmarks show SpectraLift consistently matches or exceeds leading supervised methods on key quality metrics, while outperforming state-of-the-art unsupervised baselines. Its lightweight design and robustness to spatial blur and temporal misalignment further highlight its practical utility. SpectraLift establishes a foundation for lightweight, physics-guided spectral inversion and paves the way for future advances in efficient, transparent HSI-MSI fusion models for HSI-SR. Future research will focus on enhancing the interpretability of the inverse spectral mapping process, alongside the integration of context-aware mechanisms aimed at augmenting the single-band high-resolution input scenario, while preserving the foundational advantages of the SpectaLift framework.
Acknowledgement
We thank Dr. Mario Parente for many helpful discussions on practical assessments of HSI-SR methods.